# Jet schemes, Newton polygons and continued fractions

### v2

@@ 8. Generating Series
an oriented tropical decomposition and the case $T(f_{1})\leq T(f_{t})\leq\mathbf{R}_{\geq 0}(1,1)$ can be flipped to $\mathbf{R}_{\geq 0}(1,1)\leq T(g_{1})\leq T(g_{t})$ where $g_{i}(x,y)=f_{i}(y,x)$ for all $i$.
Finally, we compute the poles of the generating series $G$ of $f$ defining a Newton plane curve singularity.
@@
Let $f$ define a Newton non-degenerate plane curve singularity. Denote by $G$ its generating series. Then $G=H+R$ (from the previous lemmas) is in $\mathbf{Z}(u,v)$ and its poles with respect to $(u,v)$ are:
$\{uv=1\}\cup\{u^{\frac{|\alpha_{l}|}{\Delta_{l}}}\cdot v^{\frac{\alpha_{l}}{\Delta_{l}}\cdot S_{l}}=e^{\frac{2ik\pi}{\Delta_{l}}}\}_{1\leq l\leq t\>\ 1\leq k\leq\Delta_{l}}$
where $\Delta_{l}=gcd(|\alpha_{l}|,\alpha_{l}\cdot S_{l})$.
@@
Recall $\alpha_{i}\cdot S_{i}=\nu_{\alpha_{i}}(f)$. $R=\frac{u}{1-uv}\sum_{i=1}^{t}r_{i}H_{T_{i}}$ where $H_{T_{i}}=\frac{u^{|\alpha_{i}|}v^{\nu_{\alpha_{i}}(f)}}{1-u^{|\alpha_{i}|}v^{\nu_{\alpha_{i}}(f)}}$. Moreover, $H=\sum_{i=1}^{t}H_{T_{i}}+\sum_{i=1}^{t+1}H_{\overset{\circ}{\sigma_{i}}}$ where for $2\leq i\leq t$, $H_{\overset{\circ}{\sigma_{i}}}=H_{T_{i}}H_{T_{i-1}}$ and for $j\in\{1,t+1\}$, write
$H_{\overset{\circ}{\sigma_{j}}}=\frac{P_{j}(u,v)}{(1-u^{|\alpha_{j}|}v^{\nu_{\alpha_{j}}(f)})}$
where $P_{j}\in\mathbf{C}[u,v]$. Thus, the set of poles $\mathcal{P}$ of $G$ verifies
$\mathcal{P}\subseteq\{uv=1\}\cup\{u^{\frac{|\alpha_{l}|}{\Delta_{l}}}\cdot v^{\frac{\alpha_{l}}{\Delta_{l}}\cdot S_{l}}=e^{\frac{2ik\pi}{\Delta_{l}}}\}_{1\leq l\leq t\>\ 1\leq k\leq\Delta_{l}}$
because we have the factorization into irreducible factors:
$u^{|\alpha_{l}|}v^{\alpha_{l}\cdot S_{l}}-1=\prod_{k=1}^{\Delta_{l}}(u^{\frac{|\alpha_{l}|}{\Delta_{l}}}\cdot v^{\frac{\alpha_{l}}{\Delta_{l}}\cdot S_{l}}-e^{\frac{2ik\pi}{\Delta_{l}}})$
  • If $T(f)=\mathbf{R}_{\geq 0}(1,1)$ with $f$ having two branches. Choose coordinates $(x,y)$ to have
    f=(yx+)(ycx+)f=(y-x+...)(y-cx+...)
@@
    $G=\frac{u^{2}v^{s}}{1-u^{2}v^{s}}+\frac{2u}{(1-uv)^{2}}\cdot\frac{u^{2}v^{2}}{1+uv}=-1+\frac{1/2}{1-uv}+\frac{1/2}{1+uv}+\frac{2u}{(1-uv)^{2}}\cdot\frac{u^{2}v^{2}}{1+uv}$
    has 1uv1-uv as a pole of multiplicity $2$, $1+uv$ of multiplicity $1$.
  • Otherwise, if also $(1,1)\notin supp(f)$, according to Observation 2, then $\nu_{\alpha}(f)>|\alpha|$ for all $\alpha\in T(f)$. Then,
@@
    with $H$ not having 1uv1-uv as a pole and $R$ having it with multiplicity $1$. Then $G$ has 1uv1-uv as a pole of multiplicity $1$. Also,
    $R=\frac{u}{1-uv}\sum_{i=1}^{t}r_{i}H_{T_{i}}$
    has $e^{\frac{2ik\pi}{\Delta_{l}}}-u^{\frac{|\alpha_{l}|}{\Delta_{l}}}\cdot v^{\frac{\alpha_{l}}{\Delta_{l}}\cdot S_{l}}$ as a pole of $R$ of multiplicity $1$. Since $G=H+R$ and $H$ doesn’t have 1uv1-uv as a pole then $G$ has $e^{\frac{2ik\pi}{\Delta_{l}}}-u^{\frac{|\alpha_{l}|}{\Delta_{l}}}\cdot v^{\frac{\alpha_{l}}{\Delta_{l}}\cdot S_{l}}$ as a pole.
@@
  • Finally if $(1,1)\in supp(f)$ then the Newton Polygon of $f$ has two faces (ie $t=2$) with $(1,1)$ as a middle vertex. Moreover, $f$ has two branches since $(1,1)\in supp(f)$ and for all $\alpha\in T(f)$, $\nu_{\alpha}(f)=|\alpha|$. So $\alpha$ and $S=(1,1)$ should correspond to poles of the form
    $\{uv=e^{\frac{2ik\pi}{|\alpha|}}\}_{1\leq k\leq|\alpha|}$
    $T(f)=\mathbf{R}_{\geq 0}\cdot\alpha\cup\mathbf{R}_{\geq 0}\cdot\beta$ where $\alpha=(a,1)$, $\beta=(1,b)$ since $(1,1)\in supp(f)$.
@@
    where $S_{1}=(0,s_{1})$ and $S_{3}=(s_{3},0)$ are the first and last vertex of the Newton polygon of $f$. This shows that the poles of $G$ are
    $\{uv=e^{\frac{2ik\pi}{|\alpha|}}\}_{1\leq k\leq|\alpha|}\cup\{uv=e^{\frac{2ik\pi}{|\beta|}}\}_{1\leq k\leq|\beta|}$

### v3

@@ 8. Generating Series
is an oriented tropical decomposition, and the case $T(f_{1})\leq T(f_{t})\leq\mathbf{R}_{\geq 0}(1,1)$ can be dealt with by flipping to $\mathbf{R}_{\geq 0}(1,1)\leq T(g_{1})\leq T(g_{t})$ where $g_{i}(x,y)=f_{i}(y,x)$ for all $i$.
Our next goal is to show that for a Newton non-degenerate $f$, $G=G(f)$ is rational and we shall compute its poles. Recall that $\frac{b}{a}$ with $\gcd(a,b)=1$ is a pole of the generating series $G$ of $f$ if we can write $G=\frac{N}{D}$ with $N,D\in\mathbf{Z}[u,v]$, $\gcd(N,D)=1$ and $1-u^{a}v^{b}$ divides $D$.
First, we start with the following lemmas which will be used in the proof of Theorem 11.
Lemma 9.
Let $G=H+R$ be a sum of $H,R\in\mathbf{Z}[u,v]$ such that $p$ is a pole of $R$ but not a pole of $H$. Then $p$ is a pole of $G$.
Proof.
Let $p=-\frac{b}{a}$ with $\gcd(a,b)=1$, and denote by $H=\frac{N_{H}}{D_{H}}$ and $R=\frac{N_{R}}{D_{R}}$ with $\gcd(N_{H},D_{H})=\gcd(N_{R},D_{R})=1$. Since $p$ is a pole of $R$ but not a pole of $H$, then $1-u^{a}v^{b}$ divides $D_{R}$ and $1-u^{a}v^{b}$ does not divide $D_{H}$. Now by writing $D_{R}=(1-u^{a}v^{b})\Delta_{R}$, then
$G=H+R=\frac{N_{H}D_{R}+N_{R}D_{H}}{(1-u^{a}v^{b})\Delta_{R}D_{H}}.$
If $p$ is not a pole of $G$, then $1-u^{a}v^{b}$ divides $N_{H}D_{R}+N_{R}D_{H}$, which in turn means that $1-u^{a}v^{b}$ divides $N_{R}D_{H}$. However, $1-u^{a}v^{b}$ is irreducible in $\mathbf{Z}[u,v]$ since $\gcd(a,b)=1$. Since $p$ is not a pole of $H$, then $\gcd(1-u^{a}v^{b},D_{H})=1$. Therefore, $1-u^{a}v^{b}$ divides $N_{R}$. So that $1-u^{a}v^{b}$ divides both $N_{R}$, $D_{R}$ which are relatively prime, which is a contradiction. Thus, the proof is complete. ∎
Lemma 10.
In $R=\mathbf{Z}[w,w^{-1}]$, the units are $R^{\times}=\pm w^{\mathbf{Z}}$.
Proof.
First note that $\pm w^{\mathbf{Z}}\subseteq R^{\times}$. Let $f\in R^{\times}$ and denote by $f^{\prime}$ its inverse in $R$. Also, let $f=\sum_{k=m}^{M}f_{k}w^{k}$ and $f^{\prime}=\sum_{i=m^{\prime}}^{M^{\prime}}f^{\prime}_{i}w^{i}$. Since $f^{\prime}f=1$, then $m+m^{\prime}=M+M^{\prime}=0$ which can be seen from comparing the least and largest exponents of $f^{\prime}f=1$. Moreover, $f$ must have one exponent $f=f_{n}w^{n}$ because otherwise there would be at least two distinct exponents in the product. As $f_{n}\in\mathbf{Z}^{\times}=\{-1,1\}$, then $f=\pm w^{n}$ with $n\in\mathbf{Z}$. This concludes our proof. ∎
Now, we end with our final theorem concerning the poles of the generating series.
@@
Let $f$ define a Newton non-degenerate plane curve singularity. Denote by $G$ its generating series. Then, $G=H+R$ (from the previous lemmas) is in $\mathbf{Z}(u,v)$ and its poles are given by:
$\{-1\}\cup\{\frac{-\alpha_{l}\cdot S_{l}}{|\alpha_{l}|}:1\leq l\leq t\}.$
@@
Recall first that $\alpha_{i}\cdot S_{i}=\nu_{\alpha_{i}}(f)$. Now $R=\frac{u}{1-uv}\sum_{i=1}^{t}r_{i}H_{T_{i}}$ where $H_{T_{i}}=\frac{u^{|\alpha_{i}|}v^{\nu_{\alpha_{i}}(f)}}{1-u^{|\alpha_{i}|}v^{\nu_{\alpha_{i}}(f)}}$, and $H=\sum_{i=1}^{t}H_{T_{i}}+\sum_{i=1}^{t+1}H_{\overset{\circ}{\sigma_{i}}}$ where for $2\leq i\leq t$, $H_{\overset{\circ}{\sigma_{i}}}=H_{T_{i}}H_{T_{i-1}}$ and for $j\in\{1,t+1\}$, set $\alpha_{t}=\alpha_{t+1}$ and write
$H_{\overset{\circ}{\sigma_{j}}}=\frac{P_{j}(u,v)}{1-u^{|\alpha_{j}|}v^{\nu_{\alpha_{j}}(f)}}$
where $P_{j}\in\mathbf{C}[u,v]$. Then, the set of poles $\mathcal{P}$ of $G$ verifies
$\mathcal{P}\subseteq\{-1\}\cup\{\frac{-\alpha_{l}\cdot S_{l}}{|\alpha_{l}|}:1\leq l\leq t\}.$
  • If $T(f)=\mathbf{R}_{\geq 0}(1,1)$ with $f$ having two branches, then choose coordinates $(x,y)$ such that
    f=(yx+)(ycx+).f=(y-x+...)(y-cx+...).
@@
    $G=\frac{u^{2}v^{s}}{1-u^{2}v^{s}}+\frac{2u}{1-uv}\cdot\frac{u^{2}v^{2}}{1-u^{2}v^{2}}=\frac{u^{2}v^{2}(1-uv+2u)}{(1-u^{2}v^{2})(1-uv)}$
    so that $G$ has 1-1 as the only pole.
  • Otherwise, if also $(1,1)\notin supp(f)$, then by Observation 2, $\nu_{\alpha}(f)>|\alpha|$ for all $\alpha\in T(f)$. Hence,
@@
    with $R$ having 1-1 as a pole. In fact,
    $R=\frac{u}{1-uv}\sum_{i=1}^{t}\frac{r_{i}\cdot u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}}{1-u^{\alpha_{i}}v^{\alpha_{i}\cdot S_{i}}}$
    which after performing computation leads to
    $R=\frac{u\sum_{i=1}^{t}r_{i}\cdot u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}\prod_{j\neq i}(1-u^{|\alpha_{j}|}v^{\alpha_{j}\cdot S_{j}})}{(1-uv)\prod_{l=1}^{t}(1-u^{|\alpha_{l}|}v^{\alpha_{l}\cdot S_{l}})}.$
    Thus, 1-1 is a pole of $R$ since 1uv1-uv does not divide its numerator. As for $I=H-H_{\overset{\circ}{\sigma}_{1}}-H_{\overset{\circ}{\sigma}_{t+1}}$, it does not have 1-1 as a pole since after a simple check we can write
    $I=\frac{q_{1}\prod_{j=2}^{t}(1-q_{j})+\sum_{i=2}^{t}q_{i}\prod_{j\neq i,i-1}(1-q_{j})}{\prod_{l=1}^{t}(1-q_{l})}$
    where $q_{j}=u^{|\alpha_{j}|}v^{\alpha_{j}\cdot S_{j}}\neq u^{k}v^{k}$ for all $k$ in the case of $I$. As for $H_{\overset{\circ}{\sigma_{1}}}$ and $H_{\overset{\circ}{\sigma_{t+1}}}$ they are of the form
    $\frac{P(u,v)}{1-u^{|\alpha|}v^{\nu_{\alpha}(f)}}$
    of course with $|\alpha|<\nu_{\alpha}(f)$ so that 1-1 is not a pole for both. This in turn means that $H$ does not have 1-1 as a pole and since $R$ has 1-1 as a pole, then $G$ has 1-1 as a pole in view of Lemma 9. As for $\frac{-\alpha_{l}\cdot S_{l}}{|\alpha_{l}|}\neq-1$ where $1\leq l\leq t$, then we can write
    $G=I+\frac{P_{1}}{1-q_{1}}+\frac{P_{t+1}}{1-q_{t+1}}+\sum_{i=1}^{t}\frac{r_{i}q_{i}u}{1-q_{i}},$
    $I=\frac{N_{I}}{\prod_{j=1}^{t}(1-q_{j})}=\frac{q_{1}}{1-q_{1}}+\sum_{i=2}^{t}\frac{q_{i}}{(1-q_{i})(1-q_{i-1})}.$
    Set $L=\{i\in[1,t+1]:\frac{\alpha_{l}\cdot S_{l}}{|\alpha_{l}|}=\frac{\alpha_{i}\cdot S_{i}}{|\alpha_{i}|}\}$. Now there exists $\frac{-b}{a}=\frac{-\alpha_{l}\cdot S_{l}}{|\alpha_{l}|}$ with $\gcd(a,b)=1$ so that we can write
    $l\in L=\{i\in[1,t+1]:1-u^{a}v^{b}\text{ divides }1-q_{i}\}=\{i_{1}<...<i_{\lambda}\}.$
    where $\lambda=\#L$. Next, we shall set
    $N=(P_{1}+q_{1})\prod_{k=2}^{t+1}(1-q_{k})+\sum_{i=2}^{t}q_{i}\prod_{k\neq i,i-1}(1-q_{k})+P_{t+1}\prod_{k=1}^{t}(1-q_{k})+\sum_{i=1}^{t}r_{i}q_{i}u\prod_{k\neq i}(1-q_{k}),$
    $D=\prod_{k=1}^{t+1}(1-q_{k})$ and keep in mind that $G=\frac{N}{D}$. Next, we consider the following cases:
    • $\blacksquare$
      if $i_{1}\neq 1$, then we write
      $D=(1-u^{a}v^{b})^{\lambda}\Delta$
      where $1-u^{a}v^{b}\nmid\Delta$. We have
      $N\equiv r_{l}u\prod_{k\neq l}(1-q_{k})+\prod_{k\neq l-1,l}(1-q_{k})+q_{l+1}\prod_{k\neq l,l+1}(1-q_{k})\mod(1-u^{a}v^{b}).$
      There exist $\varepsilon_{1},\varepsilon_{2}\in\{0,1\}$, $\Delta_{i}\neq 0\mod(1-u^{a}v^{b})$ for $i\in\{l-1,l,l+1\}$ such that,
      $N\equiv r_{l}u(1-u^{a}v^{b})^{\lambda-1}\Delta_{l}+(1-u^{a}v^{b})^{\lambda-1-\varepsilon_{1}}\Delta_{l-1}+(1-u^{a}v^{b})^{\lambda-1-\varepsilon_{1}}\Delta_{l+1}$
      modulo $1-u^{a}v^{b}$. Since $\lambda-1-\varepsilon_{i}<\lambda$, we get that $G$ has $\frac{b}{a}$ as a pole.
    • $\blacksquare$
      If $i_{1}=1$, $2\leq l\leq t$, then
      $N\equiv q_{2}\prod_{k\neq 1,2}(1-q_{k})\mod(1-u^{a}v^{b})$
      but as in the previous case, this becomes
      $N\equiv(1-u^{a}v^{b})^{\lambda-1-\varepsilon}\Delta_{1,2}\mod(1-u^{a}v^{b})$
      where $\Delta_{1,2}$ is not divisible by $1-u^{a}v^{b}$. Since $\lambda-1-\varepsilon<\lambda$, then b/a-b/a is again a pole of $G$.
    • $\blacksquare$
      If $\lambda>1$ and $i_{1}=l=1$, then
      $N\equiv q_{2}\prod_{k\neq 1,2}(1-q_{k})\equiv q_{2}(1-u^{a}v^{b})^{\lambda-2}\Delta_{2}\mod(1-u^{a}v^{b})$
      which implies as before that $G$ has b/a-b/a as a pole.
    • $\blacksquare$
      If $\lambda=i_{1}=l=1$, then
      $N\equiv(P_{1}+1+r_{1}u)\prod_{k=2}^{t+1}(1-q_{k})+q_{2}\prod_{k\neq 1,2}(1-q_{k})\mod(1-u^{a}v^{b})$
[…]
      with $1-u^{a}v^{b}\nmid\Delta^{\prime}$. Suppose by contradiction that
      $q_{2}\equiv(q_{2}-1)(P_{1}+1+r_{1}u)\mod\mathfrak{p}$
      (where $\mathfrak{p}=(1-u^{a}v^{b})$). In the quotient ring $\mathbf{Z}[u,v]/\mathfrak{p}$, we have
      $q_{2}=(q_{2}-1)(P_{1}+1+r_{1}u)$
      so that $q_{2}-1\mid q_{2}$. However, in this case
      $(q_{2}-1)=(q_{2},q_{2}-1)=\mathbf{Z}[u,v]/\mathfrak{p}=\mathbf{Z}[v,v^{-b/a}].$
      Hence, $1-q_{2}=u^{|\alpha_{2}|}v^{\alpha_{2}\cdot S_{2}}$ is a unit in the quotient ring. Moreover,
      $\mathbf{Z}[v,v^{-b/a}]=\sum_{r,s\in\mathbf{N}_{0}}\mathbf{Z}\cdot v^{\frac{1}{a}(ra-sb)}.$
      On the other hand, by Bézout’s theorem, there exist $r,s\in\mathbf{N}_{0}$ such that $ra-sb=\pm 1$.
        if rasb=1ra-sb=1 then $\mathbf{Z}[v,v^{-b/a}]=\mathbf{Z}[v^{1/a}]$ and $v^{b}$ is not invertible which is a contradiction.
        if rasb=1ra-sb=-1 then $\mathbf{Z}[v,v^{-b/a}]=\mathbf{Z}[v,v^{-1/a}]$. Next if we set $w=v^{1/a}$, then we get $u^{a}w^{ab}=1$ and
        $\mathbf{Z}[v,v^{-1/a}]=\mathbf{Z}[w^{a},w^{-1}]=\mathbf{Z}[w,w^{-1}].$
      Now $1-q_{2}$ is a unit in the ring $\mathbf{Z}[w,w^{-1}]$ so that $q_{2}=1+\varepsilon w^{n}$ with $\varepsilon=\pm 1$, $n=-b|\alpha_{2}|+a(\alpha_{2}\cdot S_{2})\in\mathbf{Z}$ by making use of Lemma 10. So in $\mathbf{Z}[u,v]/\mathfrak{p}=\mathbf{Z}[w,w^{-1}]$:
      $q_{2}+(1-q_{2})(P_{1}+1+r_{1}u)=0,\text{ hence }$
      $P_{1}+1+r_{1}u=\varepsilon w^{-n}(1+\varepsilon w^{n})=1+\varepsilon w^{-n},$
      which gives that
      $P_{1}=\varepsilon w^{-n}-r_{1}u.$
      As $u^{a}w^{ab}=1$ then $u=\varepsilon_{u}w^{-b}$ with $\varepsilon_{u}=\pm 1$. So we arrive at,
      $P_{1}=\sum_{\nu\in U_{1}}u^{|\nu|}v^{\nu\cdot S_{1}}=\sum_{\nu}(\varepsilon_{u})^{|\nu|}w^{a(\nu\cdot S_{1})-b|\nu|}$
      where $U_{1}\subset\mathbf{N}^{2}$ is a set of lattice points. Now
      $\varepsilon w^{-n}-r_{1}\varepsilon_{u}w^{-b}=P_{1}=\sum_{\nu}(\varepsilon_{u})^{|\nu|}w^{a(\nu\cdot S_{1})-b|\nu|}\in\mathbf{Z}[w,w^{-1}]$
      and $1-u^{a}v^{b}=1-\varepsilon_{u}w^{-ab+ab}=1-\varepsilon_{u}$. Then $\varepsilon_{u}=1$ because $1-u^{a}v^{b}=0\in\mathbf{Z}[w,w^{-1}]$. Next, we split the 2 cases: $P_{1}\neq 0$ and $P_{1}=0$.
      1. (1)
        if $P_{1}\neq 0$ then $\alpha_{1}\neq(1,1)$ and there exists $\beta\in U_{1}$ with
        $a(\beta\cdot S_{1})-b|\beta|=-b,$
        $\frac{\alpha_{1}\cdot S_{1}}{|\alpha_{1}|}=b/a=\frac{\beta\cdot S_{1}}{|\beta|+1}.$
        As $b=a(\beta\cdot S_{1})-b|\beta|$, then
        $a(\beta\cdot S_{1})<b|\beta|<b(1+|\beta|),$
        $b/a=\frac{\beta\cdot S_{1}}{1+|\beta|}<\frac{b}{a}\cdot\frac{|\beta|}{1+|\beta|}<b/a$
        which is clearly a contradiction.
      2. (2)
        if $P_{1}=0$ then $\varepsilon=r_{1}=1$, $b=n$ and $\alpha_{1}=(1,1)$ so that
        $uq_{2}=u(1+w^{n})=u(1+w^{b})=u(1+u^{-1})=1+u$
        in the quotient ring. Then,
        $uq_{2}=1+u-(1-u^{a}v^{b})Q,$
        for some $\mathbf{Z}[u,v]\ni Q\equiv 1\mod u$. If we write $Q=1+uU$, then
        $u^{|\alpha_{2}|}v^{\alpha_{2}\cdot S_{2}}=1-q_{2}=(1-u^{a}v^{b})U-u^{a-1}v^{b}.$
        Let us set $i=|\alpha_{2}|$ and $j=\alpha_{2}\cdot S_{2}$ in
        $u^{i}v^{j}+u^{a-1}v^{b}=w^{aj-bi}+w^{b}\neq 0.$
        Since $1-u^{a}v^{b}=0\in\mathbf{Z}[w,w^{-1}]$ and $U\in\mathbf{Z}[u,v]$ then we must have $u^{i}v^{j}+u^{a-1}v^{b}=0\in\mathbf{Z}[w,w^{-1}]$ which is not the case. Hence we arrived at a contradiction.
      Therefore, in all cases, $1-u^{a}v^{b}$ does not divide $q_{2}+(1-q_{2})(P_{1}+1+r_{1}u)$. Thus, $N\equiv(1-u^{a}v^{b})^{\lambda-1}\Delta^{\prime\prime}\mod(1-u^{a}v^{b})$ with $1-u^{a}v^{b}\nmid\Delta^{\prime\prime}$. So that $G$ has b/a-b/a as a pole.
@@
  • Finally, if $(1,1)\in supp(f)$ then the Newton Polygon of $f$ has two faces (i.e. $t=2$) with $(1,1)$ as a middle vertex. Moreover, $f$ has two branches since $(1,1)\in supp(f)$ and for all $\alpha\in T(f)$, $\nu_{\alpha}(f)=|\alpha|$. So $\alpha$ and $S=(1,1)$ should correspond to the pole $1=\frac{-\alpha\cdot S}{|\alpha|}$. Since $T(f)=\mathbf{R}_{\geq 0}\cdot\alpha\cup\mathbf{R}_{\geq 0}\cdot\beta$ where $\alpha=(a,1)$, $\beta=(1,b)$ (as $(1,1)\in supp(f)$), then
@@
    where $S_{1}=(0,s_{1})$ and $S_{3}=(s_{3},0)$ are the first and last vertex of the Newton polygon of $f$. This shows that the pole of $G$ is $1=\frac{-\alpha\cdot S}{|\alpha|}=\frac{-\beta\cdot S}{|\beta|}$ because $G\in\mathbf{Z}(u,v)\backslash\mathbf{Z}[u,v]$.